niahmkk:

what is the average rate of change of the function g(x) from n=1 to n=4? show all work to receive full credit equation: g(n) = 10(1.02)^n

Solaimon:

To find the average rate of change of the function g(n) from n = 1 to n = 4, you can use this following formula: Average Rate of Change = [g(4) - g(1)] / (4 - 1) Determine it in detail: 1. Enter 10(1.02)^4 = g(4) = 10(1.219024) to compute g(4). g(4) = 12.19024 2. Compute g(1): (1.02)^1 + g(1) For g(1), 10(1.02) 10.20 for g(1) Put these numbers into the average rate of change calculation now: [12.19024 - 10.20] / (4 - 1) is the average rate of change. [1.99024] / 3 is the average rate of change. 0.66341333333 is the average rate of change. So, the function g(n) changed at an average rate of about 0.6634 from n = 1 to n = 4. This simply indicates that as n increases from 1 to 4, the function g(n) increases by around 0.6634 units annually on average.

Stringcheese272:

idk